If you find a correlation of .50, what percent of the variance in one variable is explained by the other variable? Do not include a % sign in your answer and please round your answer to the nearest 2 decimal places. Remember, this is a percentage and a percentage is different than a proportion. A percentage is a proportion multiplied by 100.
Answer #1

Solution:- The percent of the variance in one variable is explained by the other variable is 50.

Correlation r = 0.50

Percent of variance in one variable is explained by the other variable = r * 100 = 50

Percent of variance in one variable is explained by the other variable = 0.50 * 100

Percent of variance in one variable is explained by the other variable = 50